If the displacement of a particle executing S.H.M is given by Y= 0.30sin(220t + 0.46) in meter then what is the frequency and maximum velocity of the particle?

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
On comparing it with
y = a sin (wt + phi)
Now
w = 220
But w = 2 *pi* f = 220
2 * 22*f /7 = 220
Hence
f = 35 Hz
 
Regards
Arun (askIITians forum expert)